# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

### v1

@@ Abstract
Integrating electric mobility (electric vehicles (EVs), electric trucks (ETs)) and renewable energy sources (RES) with the power grid is paramount for achieving decarbonization, efficiency, and stability. Given the rapid growth of decentralized technologies and their critical role in decarbonization, two critical challenges emerge: first, the development of a digital platform for operational coordination; and second, rigorous research into their cost–benefit profile. This paper addresses this by presenting a comprehensive cost–benefit analysis (CBA) of an AI-driven operational digital platform (ODP) designed for holistic, cross-sectoral optimization. The ODP aims to enhance energy efficiency, grid reliability, and environmental sustainability. A seven-step CBA framework, aligned with EU guidelines, quantifies economic, reliability, and environmental benefits against capital and operational expenditures, explicitly linking benefit magnitude to AI-driven ODP and optimization efficiencies, such as quantified improvements in market arbitrage from ODP, enabled forecasting, and enhanced operational efficiencies across various services. The framework is illustrated with case studies from Austria, Hungary, and Slovenia. Results show a robust benefit-cost ratio (BCR) of approximately 1.41 and a net present value (NPV) exceeding €356 million over a ten-year operational period (2026–2035). Extensive sensitivity analyses, including discount-rate variability and Monte Carlo simulations on BCR and NPV, confirm the framework’s robustness. This research underscores AI-powered integration as critical for achieving decarbonization and efficiency goals in the evolving energy and transport sectors.
@@
The European Union (EU) is undergoing a profound transformation towards sustainable energy and transport systems, characterized by the accelerated adoption of electric vehicles (EVs), electric trucks (ETs), and an increasing share of renewable energy sources (RES) in the power grid [1, 2, 3, 4]. The European Union has set the ambitious goal of achieving 60 % EV/ET sales of total vehicles by 2030 [5]. The effective integration of burgeoning EV/ET charging demands with intermittent RES generation and existing grid infrastructure presents substantial technical and economic challenges [6, 7]. These include managing the spatiotemporal mismatch between variable renewable supply and dynamic charging demand, enhancing grid flexibility to accommodate new, potentially high-impact loads, and overcoming the lack of coordinated platforms for holistic management across the energy and transport sectors. Traditional approaches often operate in silos, leading to suboptimal asset utilization, grid congestion, curtailment of renewable energy, and underutilization of emerging flexibility resources.
@@
The challenge becomes more pronounced for heavy-duty vehicles. Studies have shown that battery-electric propulsion is technically viable but economically constrained for long-haul freight applications due to battery weight, range limitations, and charging infrastructure gaps [8, 9]. Additionally, analyses comparing battery trucks to electric road systems (ERS) highlight the complexity of infrastructure investment decisions, especially in cross-border settings [3, 10]. For instance, [3] emphasizes that without coordination, smaller countries may opt out of ERS investments, even if the overall EU welfare would benefit from collaboration, underscoring the need for integrated digital solutions that can bridge national and sectoral divides.
@@
Cost-benefit analysis (CBA) offers a systematic framework for evaluating the economic, financial, and societal viability of complex integrated solutions like the one proposed herein. In European projects, standardized CBA methodologies have been developed to assess smart grid and sustainable energy projects. Notably, the European Commission Joint Research Centre (EC JRC) [11] outlines a multi-step CBA approach tailored for smart grid investments in developing countries, while [12] details a similar framework applied to a smart grid pilot in Isernia, Italy, focusing on distribution system automation. Both methodologies advocate a sequential analytical process, encompassing scope definition, benefit identification, monetization, and uncertainty assessment.
@@
However, applying these CBA frameworks to comprehensively evaluate platforms that simultaneously manage energy and transport systems, particularly those leveraging advanced AI across multiple sectors and national borders, requires careful and detailed consideration of unique benefit streams and cost structures. While existing frameworks provide a robust foundation, their application must rigorously quantify the specific contributions of AI functionalities and cross-sectoral synergies. Even recent evaluations of integrated e-mobility systems often remain confined to national boundaries or siloed sectors [13, 14]. Studies assessing the economic viability of V2G strategies and smart charging [14], or real-time optimization for EVs [15, 16], often focus on partial benefits and neglect the compounded synergies across sectors and borders that a unified, AI-driven digital platform could unlock.
@@
In this work, we propose a CBA framework to assess an AI-driven operational digital platform (ODP) that integrates EVs/ETs, RES, and the grid. While many platforms focus on specific aspects such as EV smart charging [17] or RES forecasting [18], the ODP holistically coordinates charging station availability, dynamic routing for EV/ET fleets considering energy constraints and battery health, demand response programs, and renewable energy forecasting within a unified, intelligent ecosystem. Specifically, this integrated management, orchestrated by the ODP’s AI engine is designed to boost grid flexibility by dynamically aligning EV/ET charging patterns with RES availability, improve overall system reliability through predictive load balancing and congestion management, and yield substantial economic benefits via optimized energy procurement, trading, operational cost reductions, and potentially extended EV/ET battery lifespans. To quantify these outcomes, we apply a seven-step CBA framework, thoroughly informed by [11] and [12]. Our application of this framework emphasizes the careful quantification of benefits directly attributable to the ODP’s capabilities (e.g., enhanced forecasting accuracy, superior optimization outcomes) and the synergistic effects of cross-sectoral integration, ensuring that AI-specific development, operational costs, and performance variabilities are explicitly addressed. This detailed application ensures alignment with EU best practices while providing a thorough assessment of this specific AI-driven, multi-sector and cross-border platform. We illustrate the framework to a comprehensive ODP deployments in Austria, Hungary, and Slovenia, comparing quantified benefits against investment and operational expenditures.
The main contributions of this paper are:
@@
Proposing a seven-step CBA framework to evaluate an AI-driven cross-sector ODP, explicitly identifying and monetizing AI-specific cost drivers (e.g., predictive analytics, optimization) and quantifying AI-driven benefit improvements (such as short-term RES forecasting and EV/ET charging optimization) as they propagate through the integrated transport–energy system.
@@
Demonstrating a multi-layered, AI-driven ODP architecture by enumerating its core AI modules for RES forecasting, battery-health–aware EV charging, E-VRPTW routing for electric trucks, and co-optimization of grid ancillary services, and showing how verifiable performance metrics (e.g., improvement in price forecasting, reduction in battery degradation). This helps quantify economic, reliability, and environmental benefits, while addressing large-scale computational considerations (e.g., real-time neural network inference and parallelized mixed-integer programming).
@@
Applying the proposed framework to real-world case studies in Austria, Hungary and Slovenia by utilizing country specific input data and projections (CAPEX, OPEX, RES capacity factors, EV/ET fleet sizes and their projections, and wholesale price time series) and computing present-value streams for each benefit (economic, reliability and environmental) alongside CAPEX/OPEX.
iv)
Deriving actionable policy and managerial insights (e.g., data-infrastructure requirements for 2026–2035, applicability to other EU Member States, and critical success factors such as interoperability standards and regulatory incentives for EV/ET aggregation) to inform Fit for 55 and REPowerEU objectives.
The remainder of the paper is structured as follows: Section 2 reviews related work. Section 3 presents the AI-driven ODP architecture, and the proposed CBA framework. Section 4 presents the CBA steps with a case study and application. Section 5 discusses the main results of the study and some of the limitations. Finally, Section 6 offers conclusions and future directions.
@@ 2 Related work
The pursuit of sustainable energy transitions has spurred significant research into the application of AI for optimizing the complex interplay between electric mobility, RES, and power grid operations [19, 1]. On the one hand, increasing reliance on digital platforms and smart charging programs can increase the risks of cyberattacks and natural disaster impacts on power systems [20]. The integration of EVs/ETS into the power system, which is overwhelmingly dominated by RES, has been the focus of several recent comprehensive reviews. Ref. [21] analyzes the economic dispatch and risk management challenges posed by large-scale plug-in hybrid EV/ET integration into power systems, emphasizing the need for coordinated scheduling strategies and robust optimization under uncertainty. Similarly, [22] provides a state-of-the-art review on the impact of fast EV charging on the utility sector, highlighting critical concerns such as load spikes, grid congestion, and the lack of regulatory readiness for widespread EV adoption. Flexibility needs in power systems with high levels of EV penetration and RES are reviewed by [23], who underscores the importance of adaptable system architecture to ensure reliability under dynamic conditions. On the other hand, recent literature underscores the transformative potential of AI in these domains, providing a robust academic foundation for the development and evaluation of integrated solutions like the AI-driven ODP proposed in this paper [24, 25, 26, 27]. These studies affirm that AI enhances the management of EVs, facilitates deeper RES integration, and enables more efficient and resilient grid operations, leading to considerable economic and environmental benefits. While the consensus points towards the positive impact of AI, the literature also acknowledges practical challenges such as data quality, system interoperability, cybersecurity, and the scalability of AI solutions, which are critical considerations addressed in the ODP’s design.
@@
A foundational understanding of AI’s role in the broader energy transition is provided by Ref. [28], who conducts a comprehensive review highlighting AI’s contributions to RES deployment, energy efficiency enhancements, improved grid stability through smart grid integration, and the management of demand-side dynamics, including innovative vehicle-to-grid (V2G) systems. Their findings directly support the ODP’s core objective of orchestrating EVs, ETs, RES, and grid operations through sophisticated AI-driven forecasting and optimization. Complementing this, the survey by Ref. [29] on AI in the renewable energy domain identifies key AI-based strategies, such as neural networks and advanced optimization algorithms, that bolster the integration of diverse renewable sources like solar and wind. This broad perspective informs the selection and design of the ODP’s AI components, including its Long Short-Term Memory (LSTM) models for forecasting and mixed-integer linear programming (MILP) algorithms for optimization, aiming for robust and scalable solutions.
@@
Focusing specifically on electric mobility, the systematic reviews by [30, 26, 31] on AI integration into EV energy management systems reveals significant advancements. Ref. [30] analysis of numerous studies highlights progress in energy optimization, intelligent route planning, charging demand forecasting, and predictive maintenance for EVs, primarily leveraging machine learning, deep learning, and genetic algorithms. On the distribution network level, Ref. [26] conducts an in-depth impact analysis of integrating projected EV charging loads into low-voltage systems, revealing significant concerns about voltage stability and transformer loading. Earlier work by Ref. [31] also provides a chronological review of EV/ET interactions with distribution systems, laying out both technical and regulatory gaps. The anecdotal analyses of AI techniques, by enhancing EV performance and battery health, directly underpin the ODP’s smart charging functionalities and logistics optimization modules, which notably incorporate considerations for battery degradation. The emphasis in their review on real-time adaptation to dynamic driving conditions and charging availability further informs the ODP’s AI-driven routing and intelligent charging strategies, aiming to maximize user benefits and overall system efficiency.
@@
The challenge of integrating large-scale RES into existing energy systems is another critical focus area where AI demonstrates considerable promise. Ref. [32] investigates AI’s pivotal role in managing RES within multi-energy systems, particularly for achieving carbon neutrality. They emphasize AI’s capacity for intelligent operational control, achieved through precise forecasting of RES generation and effective demand-response strategies. This aligns closely with the ODP’s objective of synchronizing EV/ET charging demand with intermittent RES availability, thereby minimizing curtailment and maximizing the utilization of clean energy. The practical applications and identified limitations discussed by [32], such as the need for sophisticated and adaptive AI methodologies, provide valuable insights that have shaped the design of the ODP’s AI engine.
@@
Parallel to advancements in RES integration and e-mobility, AI is recognized as a key enabler for the fundamental transformation of electrical grids into smarter, more resilient, and sustainable energy infrastructures. Ref. [33] explores this transformative impact, tracing the evolution from centralized to increasingly decentralized grid structures. Their work highlights AI’s crucial contributions to real-time data analytics, predictive maintenance of grid assets, and the optimization of demand-response programs. These AI-driven capabilities significantly enhance overall grid efficiency, operational flexibility, and resilience, directly corresponding to the ODP’s functions of improving grid stability and reliability through proactive, intelligent control strategies. The frameworks discussed for managing decentralized networks and ensuring interoperability inform the design of the ODP’s middleware layer and its approach to facilitating cross-border energy coordination. Such grid transformation efforts naturally lead to questions of economic justification.
@@
Consequently, the economic and environmental viability of AI-driven smart energy systems becomes a crucial aspect for their practical implementation. Ref. [34], in its analysis of the InteGRIDy project, addresses this by conducting a detailed cost and environmental benefit assessment of integrating renewable energy and smart technologies. Their study quantifies tangible economic savings and significant environmental benefits, such as reductions in CO2 emissions. This work provides a pertinent methodological precedent and empirical support for the comprehensive seven-step CBA framework employed in our evaluation of the ODP, particularly in its approach to quantifying diverse benefits like reduced operational costs, enhanced market revenues, and monetized emission reductions across multiple national contexts.
@@
Overall, this body of research underscores the significant potential of AI to revolutionize energy and transport systems. While individual studies demonstrate clear benefits in specific domains such as EV management, RES forecasting, or grid optimization, a clear opportunity remains for holistic platforms, like the proposed ODP, that integrate these functionalities to achieve synergistic, system-wide gains. The existing literature thus provides a strong impetus for the development and rigorous evaluation of such comprehensive AI-driven solutions. This study distinguishes itself by conducting a comprehensive CBA of a holistic, AI-driven ODP spanning multiple countries and sectors, with a particular focus on explicitly quantifying the incremental benefits derived from the integrated AI functionalities, an aspect often treated qualitatively or with a narrower scope in prior work. Our work on the CBA of ODP that integrates EVs/ETS, RES and grid responds to this by not only proposing such an integrated platform but also by conducting a detailed cost-benefit analysis to quantify its overall value proposition, directly addressing practical implementation challenges related to data management, AI model integration, and scalability, which are key considerations in our design.

### v2

@@ Abstract
Integrating electric mobility, such as electric vehicles (EVs), electric trucks (ETs), and renewable energy sources (RES) with the power grid is paramount for decarbonization, efficiency, and stability. Two gaps remain: the need for digital platforms that coordinate operations across sectors and borders, and the lack of a rigorous framework on their cost–benefit profile. This paper presents a comprehensive cost–benefit analysis (CBA) of an AI-driven operational digital platform (ODP) designed for holistic, cross-sectoral and cross-border optimization. The ODP targets improvements in energy efficiency, grid reliability, and environmental sustainability. We develop a seven-step CBA framework that links each benefit stream to the platform’s architecture and quantifies incremental gains relative to the existing system, while accounting for investment and operating costs. The framework is demonstrated with case studies for Austria, Hungary, and Slovenia. Results indicate a benefit-cost ratio (BCR) of about 1.41 and a net present value (NPV) exceeding €356 million over 2026–2035. Robustness is examined through extensive sensitivity analyses that vary discount rates, cost components, and adoption trajectories, as well as through Monte Carlo simulations that capture uncertainty in BCR, NPV, data availability, and AI accuracy. The findings support the viability of investing in digital platforms for cross-sectoral and cross-border integration, and highlight the role of ODPs in advancing decarbonization and efficiency in the mobility–power nexus.
@@
The European Union is accelerating the shift to sustainable energy and transport, with rising adoption of electric vehicles (EVs) and electric trucks (ETs) and a growing share of renewable energy sources (RES) in the grid [Motlagh et al., 2025, International Energy Agency (2021), IEA, Börjesson and Proost, 2025, Breed et al., 2021]. The EU targets EVs/ETs to reach 60% of new vehicle sales by 2030 [Mądziel and Campisi, 2023]. The effective integration of burgeoning EV/ET charging demands with intermittent RES generation and existing grid infrastructure presents substantial technical and economic challenges [Yuvaraj et al., 2024, Psarros and Papathanassiou, 2023]. These include managing the spatiotemporal mismatch between variable renewable supply and dynamic charging demand, enhancing grid flexibility to accommodate new, potentially high-impact loads, and overcoming the lack of coordinated platforms for holistic management across the energy and transport sectors. Traditional approaches often operate in silos, leading to suboptimal asset utilization, grid congestion, curtailment of renewable energy, and underutilization of emerging flexibility resources.
@@
The challenge becomes more pronounced for heavy-duty vehicles. Studies have shown that battery-electric propulsion is technically viable but economically constrained for long-haul freight applications due to battery weight, range limitations, and gaps in charging infrastructure [Çabukoglu et al., 2018, Cunanan et al., 2021]. Additionally, analyses comparing battery trucks to electric road systems (ERS) highlight the complexity of infrastructure investment decisions, especially in cross-border settings [Börjesson and Proost, 2025, Deshpande et al., 2023]. For instance, Börjesson and Proost [2025] emphasizes that without coordination, smaller countries may opt out of ERS investments, even if the overall EU welfare would benefit from collaboration, underscoring the need for integrated digital solutions that can bridge national and sectoral divides.
@@
Cost-benefit analysis (CBA) provides a systematic framework for evaluating the economic, financial, and societal viability of complex integrated solutions like the one proposed herein. In European projects, standardized CBA methodologies have been developed to assess smart grid and sustainable energy projects. Notably, the European Commission Joint Research Centre (EC JRC) [EU Commission JRC, 2012] outlines a multi-step CBA approach tailored for smart grid investments in developing countries. On the other hand, while Stromsather et al. [2018] details a similar framework applied to a smart grid pilot project in Isernia, Italy, focusing on distribution system automation. Both methodologies advocate a sequential analytical process, encompassing scope definition, benefit identification, monetization, and uncertainty assessment.
@@
Applying these CBA frameworks to comprehensively evaluate smart grid solutions that simultaneously manage energy and transport systems, particularly those leveraging high RES penetration, requires careful and detailed consideration of unique benefit streams and cost structures. While existing frameworks provide a robust foundation for smart grid solutions, there is no CBA framework for AI-driven operational digital platforms (ODPs) that rigorously quantify the specific contributions of AI functionalities and cross-sectoral and cross-border synergies. Even recent evaluations of integrated e-mobility systems often remain confined to national boundaries or siloed sectors [de Saxe et al., 2023, Greaker et al., 2022]. Studies assessing the economic viability of V2G strategies and smart charging [Greaker et al., 2022], or real-time optimization for EVs [Feng and Figliozzi, 2013, Vázquez-Canteli and Nagy, 2019], often focus on partial benefits and neglect the compounded synergies across sectors and borders that a unified, AI-driven digital platform could unlock. There is currently no CBA framework that utilizes modern AI and digital platforms to achieve cross-sector and cross-border integration and scalable smart technologies.
@@
In this work, we propose a CBA framework to assess an ODP that integrates EVs/ETs, RES, and the grid. While many platforms focus on specific functions such as EV smart charging [Schott et al., 2019] or RES forecasting [Ahmed et al., 2020], the ODP in this work takes a broader view. It coordinates charging-station availability, dynamic routing for EV/ET fleets with energy and battery-health constraints, demand response programs, and renewable forecasting within one intelligent ecosystem. Specifically, this integrated management, orchestrated by the ODP’s AI engine, is designed to boost grid flexibility by dynamically aligning EV/ET charging patterns with RES availability. This, in turn, enhances overall system reliability through predictive load balancing and congestion management, yielding substantial economic benefits via optimized energy procurement, trading, and operational cost reductions, as well as potentially extending EV/ET battery lifespans. To quantify these outcomes, we apply a seven-step CBA framework, thoroughly informed by EU Commission JRC [2012] and Stromsather et al. [2018]. Although we follow EU guidelines for developing our CBA framework, our contribution extends beyond application to the development of a stylized, holistic methodology that explicitly considers the ODP core architecture. In this regard, there is no prior CBA framework that can be applied to operational digital platforms, given the significant investments required for integrating digital platforms and AI models. Additionally, we tested the framework with a practical case study that considered the interests of stakeholders qualitatively and the financial and economic feasibility quantitatively. Our application of this framework emphasizes the careful quantification of benefits directly attributable to the ODP’s capabilities and the synergistic effects of cross-sectoral integration, ensuring that AI-specific development, operational costs, and performance variabilities are explicitly addressed. Regarding the mobility dimension of the ODP, it is important to note that the proposed CBA framework is not an operational optimization of EV/ET logistics. Rather, it quantifies the incremental benefits of the AI-driven ODP by parameterizing EV/ET characteristics together with RES and grid interactions. Key factors include higher energy use per kilometer, reduced charging flexibility due to duty schedules. These parameters are incorporated into the benefit and cost channels (e.g., curtailment reduction, peak demand management, battery degradation savings), allowing EV/ET-specific constraints to influence system-level outcomes without requiring a full routing or scheduling model. This detailed application ensures alignment with EU practices while providing an investment decision-making CBA framework for AI-driven, multi-sector, and cross-border platform. Benefits are anchored to the ODP-enabled system compared with a system without ODP, using it as a baseline. Then, costs and benefits are quantified through platform-to-KPI linkages (e.g., curtailment reduction, peak shifting, imbalance reduction, congestion reduction, and reliability improvements). Environmental benefits valuation uses country-specific, time-resolved factors rather than static averages, and uncertainty analysis explicitly tests AI/data performance and adoption trajectories. This integrated, cross-sector, and cross-border framework offers a methodological advance beyond standard CBAs, which focus on single infrastructures, and yields a novel model for evaluating digital platforms that coordinate mobility, energy, and grid systems. We illustrate the CBA framework using a deployment of the ODP in Austria, Hungary, and Slovenia. Incorporating country-specific, technology, and economic assumptions, the case study reflects the cross-sector and cross-border nature of the ODP as it integrates diverse EV, ET, RES, and grid conditions. Therefore, the main contributions of the paper are threefold:
@@
Proposing a seven-step CBA framework to evaluate an AI-driven cross-sector, and cross-border ODP, explicitly identifying and monetizing AI-specific cost drivers and quantifying AI-driven benefit improvements (such as short-term RES forecasting and EV/ET charging optimization) as they propagate through the integrated transport–energy system.
@@
Demonstrating a multi-layered, AI-driven ODP architecture by enumerating its core AI modules for RES forecasting, battery-health–aware EV charging, E-VRPTW routing for electric trucks, and co-optimization of grid ancillary services, and showing how verifiable performance metrics. This helps quantify economic, reliability, and environmental benefits, and the interconnectedness of each layer in its architectural core.
@@
Applying the proposed framework to real-world case studies in Austria, Hungary, and Slovenia by utilizing country-specific input data and projections (CAPEX, OPEX, RES capacity factors, EV/ET fleet sizes and their projections, and wholesale price time series) and computing present-value streams for each benefit alongside CAPEX/OPEX. This gives actionable policy and managerial insights to inform Fit for 55 and REPowerEU objectives.
The remainder of the paper is structured as follows: Section 2 reviews related work. Section 3 presents the AI-driven ODP architecture and the proposed CBA framework. Section 4 presents the CBA steps with a case study and application. Section 5 discusses the main results of the study and some of the limitations. Finally, Section 6 offers conclusions and future directions.
@@ 2 Related work
The pursuit of sustainable energy transitions has spurred significant research into the application of AI for optimizing the complex interplay between electric mobility, RES, and power grid operations [Huang et al., 2025, Motlagh et al., 2025]. On the one hand, increasing reliance on digital platforms and smart charging programs can increase the risks of cyberattacks and the impact of natural disaster on power systems [Diaba et al., 2024]. The integration of EVs/ETs into the power system, which is overwhelmingly dominated by RES, has been the focus of several recent comprehensive reviews. Ref. Peng et al. [2012] analyzes the economic dispatch and risk management challenges posed by large-scale plug-in hybrid EV/ET integration into power systems, emphasizing the need for coordinated scheduling strategies and robust optimization under uncertainty. Similarly, Shariff et al. [2022] provides a state-of-the-art review on the impact of fast EV charging on the utility sector, highlighting critical concerns such as load spikes, grid congestion, and the lack of regulatory readiness for widespread EV adoption. Flexibility needs in power systems with high levels of EV penetration and RES are reviewed by Kaushik et al. [2022], who underscores the importance of adaptable system architecture to ensure reliability under dynamic conditions. On the other hand, recent literature highlights the transformative potential of AI in these domains, providing a robust academic foundation for the development and evaluation of integrated solutions, such as the AI-driven ODP proposed in this paper [Zhang et al., 2025, Child and Breyer, 2017, Rahman et al., 2022, Kumar et al., 2019]. These studies affirm that AI enhances the management of EVs, facilitates deeper RES integration, and enables more efficient and resilient grid operations, leading to considerable economic and environmental benefits. While the consensus points towards the positive impact of AI, the literature also acknowledges practical challenges such as data quality, system interoperability, cybersecurity, and the scalability of AI solutions, which are critical considerations addressed in the ODP’s design.
@@
A foundational overview of AI’s role in the energy transition is provided by Ref. Wang et al. [2025]. The review highlights AI’s contributions to RES deployment and energy-efficiency improvements. It also covers enhanced grid stability via smart-grid integration and the management of demand-side dynamics, including vehicle-to-grid (V2G) systems. Their findings directly support the ODP’s core objective of orchestrating EVs, ETs, RES, and grid operations through sophisticated AI-driven forecasting and optimization. Complementing this, the survey by Ref. Al Nasim et al. [2024] on AI in the renewable energy domain identifies key AI-based strategies, such as neural networks and advanced optimization algorithms, that enhance the integration of diverse renewable energy sources, such as solar and wind. This broad perspective informs the selection and design of the ODP’s AI components, including its Long Short-Term Memory (LSTM) models for forecasting and mixed-integer linear programming (MILP) algorithms for optimization, with the goal of achieving robust and scalable solutions.
@@
Focusing specifically on electric mobility, the systematic reviews by Arévalo et al. [2024], Rahman et al. [2022], Arias-Londoño et al. [2020] on the integration of AI into EV energy management systems reveal significant advancements. Ref. Arévalo et al. [2024] analyzes numerous studies that highlight progress in energy optimization, intelligent route planning, charging demand forecasting, and predictive maintenance for EVs, primarily leveraging machine learning, deep learning, and genetic algorithms. On the distribution network level, Ref. Rahman et al. [2022] conducts an in-depth impact analysis of integrating projected EV charging loads into low-voltage systems, revealing significant concerns about voltage stability and transformer loading. Earlier work by Ref. Arias-Londoño et al. [2020] also provides a chronological review of EV/ET interactions with distribution systems, laying out both technical and regulatory gaps. The anecdotal analyses of AI techniques, by enhancing EV performance and battery health, directly underpin the ODP’s smart charging functionalities and logistics optimization modules, which notably incorporate considerations for battery degradation.
@@
The challenge of integrating large-scale RES into existing energy systems is another critical focus area where AI demonstrates considerable promise. Ref. Liu et al. [2022] investigates AI’s pivotal role in managing RES within multi-energy systems, particularly for achieving carbon neutrality. They emphasize AI’s capacity for intelligent operational control, achieved through precise forecasting of RES generation and effective demand-response strategies. This aligns closely with the ODP’s objective of synchronizing EV/ET charging demand with intermittent RES availability, thereby minimizing curtailment and maximizing the utilization of clean energy. The practical applications and identified limitations discussed by Liu et al. [2022], such as the need for sophisticated and adaptive AI methodologies, provide valuable insights that have shaped the design of the ODP’s AI engine.
@@
Parallel to advancements in RES integration and e-mobility, AI is recognized as a key enabler for the fundamental transformation of electrical grids into smarter, more resilient, and sustainable energy infrastructures. Ref. Rajaperumal and Columbus [2025] explores this transformative impact, tracing the evolution from centralized to increasingly decentralized grid structures. Their work highlights AI’s crucial contributions to real-time data analytics, predictive maintenance of grid assets, and the optimization of demand-response programs. These AI-driven capabilities significantly enhance overall grid efficiency, operational flexibility, and resilience. This directly corresponds to the ODP’s functions of improving grid stability and reliability through proactive, intelligent control strategies. The frameworks discussed for managing decentralized networks and ensuring interoperability inform the design of the ODP’s middleware layer and its approach to facilitating cross-border energy coordination. Such grid transformation efforts naturally lead to questions of economic justification.
@@
Consequently, the economic and environmental viability of AI-driven smart energy systems becomes a crucial aspect for their practical implementation. Ref. Gudlaugsson et al. [2023] addresses this issue in its analysis of the InteGRIDy project, by conducting a detailed cost and environmental benefit assessment of integrating renewable energy and smart technologies. Their study quantifies tangible economic savings and significant environmental benefits, such as reductions in CO2 emissions. This work provides a pertinent methodological precedent and empirical support for the comprehensive seven-step CBA framework employed in our evaluation of the ODP.
@@
This body of research highlights the significant potential of AI to revolutionize energy and transport systems. Despite individual studies demonstrate clear benefits in specific domains such as EV management, RES forecasting, or grid optimization, a clear research gap remains for a holistic digital platforms that integrate mobility and energy sectors. This study, therefore, proposes a novel CBA framework employing a comprehensive AI-driven ODP architecture for a cross-sector and cross-border energy-transport efficiency. The main objective is quantifying the incremental costs and benefits derived from the integrated AI functionalities, an aspect often treated qualitatively or with a narrower scope in prior work. The proposed CBA framework not only integrates EVs/ETs, RES, and grid, but also conducts a detailed CBA to quantify the potential benefits and expenditures.